OpenStudy (hannahduhe):

the vertex angle of an isosceles triangle measure eight times the measure of a base angle. Find the measure of a base angle. A.18 B.24 C.36 D.43

OpenStudy (mathmale):

Didn't we work on this very same problem a minute or two back? Or was that another O. S. user?

OpenStudy (hannahduhe):

i acidentlly closed it

OpenStudy (mathmale):

Try your utmost to remember what was said in that post. You'll need to set up an equation in the variable x. Start by representing each of the 2 base angles by x. What is an appropriate representation of the vertex angle if the vertex angle is 8 times the measure of one of the base angles? Write your equation. these 3 angles add up to what?

OpenStudy (hannahduhe):

so the equation would be 1/2 times base times height

OpenStudy (mathmale):

No: that's the formula for the area of a triangle. Maybe you should draw the triangle in question. Note that it's isosceles. What does that mean?

OpenStudy (hannahduhe):

an isosceles triangle has two equal sides right

OpenStudy (mathmale):

Right, and so the base angles are equal. Represent each of the base angles with a letter. What'll it be?

OpenStudy (hannahduhe):

x

OpenStudy (mathmale):

Great. What represents the measure of the other base angle?

OpenStudy (hannahduhe):

vertex

OpenStudy (mathmale):

Yes, but that vertex angle is measured (e. g., in degrees or in radians). We have to represent it algebraically. How large does the problem statement say this angle is?

OpenStudy (hannahduhe):

8 times the measure

OpenStudy (mathmale):

Right. 8 times the measure of ... what?

OpenStudy (hannahduhe):

the base angle

OpenStudy (mathmale):

One of the base angles. if one base angle is represented by x, what is 8 times that?

OpenStudy (hannahduhe):

8

OpenStudy (mathmale):

8 times what? The measure of one of the base angles is x. The measure of the vertex angle is 8 times that. Represent the measure of the vertex angle here, please.

OpenStudy (hannahduhe):

8x

OpenStudy (mathmale):

Very good. Now, you have three angles inside this triangle. They are x, x and 8x. Add those together and set your sum equal to ... what value?

OpenStudy (hannahduhe):

2x+8x=180?

OpenStudy (mathmale):

Yes, and 2x + 8x = ?

OpenStudy (hannahduhe):

10x?

OpenStudy (mathmale):

Yes, and so 10x = ?

OpenStudy (hannahduhe):

18?

OpenStudy (mathmale):

Yes, and now we have to check that. We know that x plus x plus 8x should equal 180 deg. Subst. 18 for x in that statement. Does it hold true, or not?

OpenStudy (hannahduhe):

18+18+8*18=180

OpenStudy (mathmale):

36+8*18 = ?

OpenStudy (hannahduhe):

180

OpenStudy (mathmale):

And so x=18 degrees is correct!

OpenStudy (hannahduhe):

yes thank u for ur time my teacher doents teach she shows us videos

OpenStudy (mathmale):

Happy to be of help, Hannah! Bye for now!

OpenStudy (hannahduhe):

bye